OpenStudy (anonymous):

Which of the following best describes using a slope of -5/10 on a graph? A. Move left 5 and up 10. B. Move down 5 and left 10. C. Move down 5 and right 10. D. Move up 5 and right 10.

OpenStudy (anonymous):

slope = rise/run = y values/x values = -5/10 = move down five/ move 10 to the right

OpenStudy (anonymous):

the answers C :)

OpenStudy (jhannybean):

slope = rise/run = -5/10 can you use this to solve? :)

OpenStudy (anonymous):

alternately if your slope is -5/10, the fraction can be rewritten as 5/-10 where you could go 5 units up and 10 units to the left :)

OpenStudy (anonymous):

Oh ok, I see. Thank you very much for your help! :) @sizzy84

OpenStudy (jhannybean):



OpenStudy (anonymous):

Wow, this makes it really clear. Down 5, 10 right. So C.

OpenStudy (jhannybean):

mmhmm.

OpenStudy (anonymous):

Thank you! @Jhannybean

OpenStudy (jhannybean):

np